OpenStudy (anonymous):

Solve for y: 2 over 3y + 2 = 8

OpenStudy (anonymous):

\[ \frac 2{3y+2}=8 \]

OpenStudy (anonymous):

The first step is multiplication.

OpenStudy (owlcoffee):

\[\frac{ 2 }{ 3y+2 }=8\] it's a proportion, so you can multiply like this: \[2=8(3y+2)\] You can continue from there on?